OpenStudy (anonymous):

The population of a type of local dragonfly can be found using an infinite geometric series where a1 = 65 and the common ratio is 1/6. Find the sum of this infinite series that will be the upper limit of this population. A. 78 B. 28 C. 11 D.32

OpenStudy (sdfgsdfgs):

do u know the general form of a geometric series with a1 and the common ratio?

OpenStudy (sdfgsdfgs):

an = a1(CR)^n

OpenStudy (anonymous):

What is the C?

OpenStudy (sdfgsdfgs):

why u said the ans is C?

OpenStudy (anonymous):

No i mean you put in the formula the letter c

OpenStudy (sdfgsdfgs):

Oooo sorry i mean to say an = a1*(Common Ratio)^n

OpenStudy (anonymous):

Oh okay . What woud the 'an' be though a1= 65 and cr = 1/6 How would you find an or letter n

OpenStudy (sdfgsdfgs):

Geometric series is a series of numbers... 1st no. = a1 2nd no. = a2 = a1*CR 3rd no. = a3 = a2*CR=a1*CR*CR .... nth no.=an=a1*(CR)^n

OpenStudy (sdfgsdfgs):

U may need to read up on geometric series before u can do this prob. plz read ur study material or check out wikipedia: https://en.wikipedia.org/wiki/Geometric_series

OpenStudy (anonymous):

@sdfgsdfgs

OpenStudy (sdfgsdfgs):

yes @Lala3712 so do u understand what a geometric series is now?

OpenStudy (anonymous):

Yes

OpenStudy (sdfgsdfgs):

good :) see if u can understand this.. http://openstudy.com/study#/updates/536fda91e4b09cc1503c7c34